INAF queue observing, October - November 2014

Observer: E. Sani, R. Carini
Telescope Operator: D. Gonzalez - Huerta, S. Allandon
Support Astronomer: M. Edwards, D. Thompson (on call)
Instrument: MODS, PARALLEL, LBC bin, LUCI

October 26-27

At the beginning of the night humidity is 97% and temperature close to the dewpoint. The enclosure stays close.
Night time: 11h
Bad weather: 11h

October 27-28

We start observations at 1:45 UT<br />
At the beginning of observations we have thin cirrus above the orison (NO direction), low wind ans humidity at 26%.
N.B. The seeing from the GCS is overestimated by 10% because the MODS guider is out of focus.
During night humidity is <10%, no wind, clear sky, variable seeing: 0.8 - 1.1 in red
Technical problems with modsDisp.
PARALLEL work failed. Preset telescope missing in the OBs (new, unknown).
Night time: 10h
Science-time: 9h
Tech problems: 40m
Switch instrument: 20m

Project_RUN Objects Filter ExpTime Notes
Standard Feige110 dual grating + red grating - airmass 1.4. Default 5" slit
Fusi Pecci Q B517 dual grating 45m COMPLETED
Fusi Pecci R H2 dual grating 45m COMPLETED
Saracco F XLLSM1 red grating 2.30 h parallel mode failed
Dotto KC46 V- ,R -Bessel 1 h  
Clementini IZw 18 V- I- Bessel 10m  

October 28-29

When the Dome is opened the humidity is 76%, no wind.
Variable seeing at the beginning of the scientific operations.
ModsDisp works well.
Colimation and copointing problems
Night time: 10h 45m
Science-time: 8h15m
Tech problems: 2h 15m
Switch instrument: 15m

Project_RUN Objects Filter ExpTime Notes
Standard feige110 dual and red grating   airmass 1.5
Fusi Pecci S H7 dual grating 1h COMPLETED
Fusi Pecci T G002 dual grating 45 m COMPLETED
Saracco F XLLSM1 red grating + LBCred 2 h 45 m  
Israel A1 catsj021950 dual grating 15 m  
Standard star SA98   airmass 1.23

October 29-30

When the Dome is opened the humidity is 8%, no wind, the sky is clear, seeing 0.6
During Cresci integration we need to restart the GCS.
LBCred one or more systems failed.
Night time: 11h
Science-time: 10h
Tech problems: 35m
Switch instrument: 20m

Project_RUN Objects Filter ExpTime Notes
Cresci DSF2237b Ks 2h PROGRAM COMPLETED
Standard feige110 dual and red grating   airmass 1.5
Saracco F XLLSM1 red grating + LBC red 2 h 30 m COMPLETED
Dotto KC46 V- ,R -Bessel 1 h 15m PROGRAM COMPLETED
Clementini IZw 18 V- I- Bessel 10m  

October 30 - 31

When the Dome is opened the sky is partly cloudy, the humidity 40%, velocity of wind is 10 m/s After the transit thick clouds
Night time: 11h
Science-time: 9h
Bad weather: 1h 30m
Switch instrument: 20m

Project_RUN Objects Filter ExpTime Notes
Nascimbeni HAT32 dual grating + red channel (LBC) 5 h 30m COMPLETED
standard g191-b2b dual grating    
Antoniucci dtauri dual grating + imaging   to be repeated
Antoniucci dtauri Ks   to be repeated

October 31 - November 01

House keeping problems for LBCBlue. Fail saving mode that stops observations. Is anyhow offered for science
00:36 is not possible open the dome, the temperature is too near the dew point and the humidity is 99.9%, it's raining on the horizon.
04.35 dome closed, strong wind, the temperature reaches the dew point.
09:45 dwepoint drops down, but still wind gusts are above 20 m/s.
Night time: 11h
Bad weather 11h

November 01 - November 02

At the beginning of the night the TO is waiting a wind decrease to open the dome.
Since we are closed we live the beginning of the night to the AO technicians.
After the first targets, the TO closes the dome due to high humidity and strong wind.
We stay closed the rest of the night. Night time: 11h
Science-time: 2h
Bad weather: 9h

Project_RUN Objects Filter ExpTime Notes
standard HIP117774 150+Ks    
Piconcelli_12_A7 2346-0016 150+Ks 45m COMPLETED? passing clouds
-- %USERSIG{EleonoraSani - 2014-11-02}%

Comments

 
Topic revision: r2 - 20 Dec 2014, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback